OpenStudy (anonymous):

I NEED HELP WILL MEDAL hoose the equation below that represents the line that passes through the point (−2, −1) and has a slope of 5. y − 1 = 5(x − 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y − 2 = 5(x − 1)

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)